# Structure-preserving contrastive learning for spatial time series

### v3

@@ Acknowledgments
Appendix A Theoretical optimal values of the losses
For ${\mathcal{L}}_{\text{TS2Vec}}$, a value of 0 is reached when ${\bm{z}}_{i,t}^{\prime}$ and ${\bm{z}}_{i,t}^{\operatorname*{{\prime\prime}}}$ are identical. Similarly, the optimal case of ${\mathcal{L}}_{\text{SoftCLT}}$ is when the samples with soft assignments close to 1 are identical, while dissimilar samples have soft assignments close to 0. The topology-preserving loss ${\mathcal{L}}_{\text{Topo}}$ is 0 when the topologically relevant distances remain the same in the latent space as in the original space, i.e., ${\bm{A}}^{X}\left[\bm{\pi}^{X}\right]={\bm{A}}^{Z}\left[\bm{\pi}^{X}\right]$ and ${\bm{A}}^{X}\left[\bm{\pi}^{Z}\right]={\bm{A}}^{Z}\left[\bm{\pi}^{Z}\right]$. Finally, ${\mathcal{L}}_{\text{GGeo}}$ approximates the distortion measure of isometry and is ideally 0, but can be negative when $\text{Tr}(\tilde{H}_{i})<2$, as the approximation of $\tilde{H}_{i}$ is kernel-based depending on a hyperparameter $h$.
Appendix B Detailed results on UEA datasets
This section provides detailed comparisons of evaluation results for the used 28 datasets in the UEA archive. Tables B.1 and B.2 present the results of classification accuracy. Tables B.3 and B.4 present the training time for self-supervised representation learning.
Table B.1: Detailed evaluation of classification accuracy on spatial datasets in the UEA archive.
@@ Appendix B Detailed results on UEA datasets
Table B.2: Detailed evaluation of classification accuracy on non-spatial datasets in the UEA archive.
@@
Table B.3: Detailed representation training time per epoch (unit: s) on spatial datasets in the UEA archive.
@@
ArticularyWordRecognition 3.799 (1.00$\times$) 5.61 (1.48$\times$) 5.863 (1.54$\times$) 3.772 (0.99$\times$) 5.77 (1.52$\times$) 5.983 (1.57$\times$)
BasicMotions 0.475 (1.00$\times$) 0.685 (1.44$\times$) 0.709 (1.49$\times$) 0.457 (0.96$\times$) 0.687 (1.45$\times$) 0.711 (1.50$\times$)
CharacterTrajectories 20.640 (1.00$\times$) 30.863 (1.50$\times$) 33.32 (1.61$\times$) 20.652 (1.00$\times$) 30.948 (1.50$\times$) 33.18 (1.61$\times$)
Cricket 1.903 (1.00$\times$) 2.653 (1.39$\times$) 5.437 (2.86$\times$) 1.904 (1.00$\times$) 2.655 (1.40$\times$) 5.436 (2.86$\times$)
ERing 0.319 (1.00$\times$) 0.482 (1.51$\times$) 0.487 (1.53$\times$) 0.316 (0.99$\times$) 0.483 (1.51$\times$) 0.49 (1.54$\times$)
EigenWorms 19.862 (1.00$\times$) 23.823 (1.20$\times$) 149.05 (7.50$\times$) 20.224 (1.02$\times$) 24.856 (1.25$\times$) 150.7 (7.59$\times$)
Epilepsy 1.737 (1.00$\times$) 2.49 (1.43$\times$) 2.753 (1.58$\times$) 1.686 (0.97$\times$) 2.506 (1.44$\times$) 2.755 (1.59$\times$)
Handwriting 1.875 (1.00$\times$) 2.771 (1.48$\times$) 2.959 (1.58$\times$) 1.88 (1.00$\times$) 2.775 (1.48$\times$) 2.987 (1.59$\times$)
LSST 29.786 (1.00$\times$) 45.273 (1.52$\times$) 45.162 (1.52$\times$) 29.859 (1.00$\times$) 45.216 (1.52$\times$) 45.154 (1.52$\times$)
Libras 2.081 (1.00$\times$) 3.142 (1.51$\times$) 3.142 (1.51$\times$) 2.085 (1.00$\times$) 3.135 (1.51$\times$) 3.141 (1.51$\times$)
NATOPS 1.953 (1.00$\times$) 2.989 (1.53$\times$) 2.949 (1.51$\times$) 2.085 (1.07$\times$) 3.147 (1.61$\times$) 3.159 (1.62$\times$)
PEMS-SF 3.413 (1.00$\times$) 5.069 (1.49$\times$) 5.38 (1.58$\times$) 3.415 (1.00$\times$) 5.064 (1.48$\times$) 5.399 (1.58$\times$)
RacketSports 1.781 (1.00$\times$) 2.685 (1.51$\times$) 2.664 (1.50$\times$) 1.771 (0.99$\times$) 2.711 (1.52$\times$) 2.665 (1.50$\times$)
UWaveGestureLibrary 1.699 (1.00$\times$) 2.395 (1.41$\times$) 2.788 (1.64$\times$) 1.776 (1.05$\times$) 2.595 (1.53$\times$) 2.99 (1.76$\times$)
Avg. over spatial datasets 6.523 sec/epoch 1.46$\times$ 2.12$\times$ 1.00$\times$ 1.48$\times$ 2.15$\times$
Table B.4: Detailed representation training time per epoch (unit: s) on non-spatial datasets in the UEA archive.
@@
AtrialFibrillation 0.182 (1.00$\times$) 0.258 (1.42$\times$) 0.369 (2.03$\times$) 0.177 (0.97$\times$) 0.259 (1.42$\times$) 0.366 (2.01$\times$)
DuckDuckGeese 0.617 (1.00$\times$) 0.973 (1.58$\times$) 1.059 (1.72$\times$) 0.621 (1.01$\times$) 0.968 (1.57$\times$) 1.104 (1.79$\times$)
EthanolConcentration 4.939 (1.00$\times$) 6.655 (1.35$\times$) 20.128 (4.08$\times$) 4.89 (0.99$\times$) 6.664 (1.35$\times$) 20.182 (4.09$\times$)
FaceDetection 70.709 (1.00$\times$) 109.6 (1.55$\times$) 108.83 (1.54$\times$) 71.104 (1.01$\times$) 107.523 (1.52$\times$) 107.092 (1.51$\times$)
FingerMovements 3.826 (1.00$\times$) 5.67 (1.48$\times$) 5.706 (1.49$\times$) 3.779 (0.99$\times$) 5.671 (1.48$\times$) 5.716 (1.49$\times$)
HandMovementDirection 2.221 (1.00$\times$) 3.353 (1.51$\times$) 4.151 (1.87$\times$) 2.226 (1.00$\times$) 3.334 (1.50$\times$) 4.142 (1.86$\times$)
Heartbeat 2.811 (1.00$\times$) 4.218 (1.50$\times$) 5.22 (1.86$\times$) 2.818 (1.00$\times$) 4.216 (1.50$\times$) 5.221 (1.86$\times$)
JapaneseVowels 3.211 (1.00$\times$) 4.871 (1.52$\times$) 4.821 (1.50$\times$) 3.199 (1.00$\times$) 4.846 (1.51$\times$) 4.83 (1.50$\times$)
MotorImagery 7.450 (1.00$\times$) 9.637 (1.29$\times$) 51.0 (6.85$\times$) 7.475 (1.00$\times$) 9.659 (1.30$\times$) 50.881 (6.83$\times$)
PhonemeSpectra 42.956 (1.00$\times$) 63.801 (1.49$\times$) 70.578 (1.64$\times$) 43.015 (1.00$\times$) 63.807 (1.49$\times$) 70.806 (1.65$\times$)
SelfRegulationSCP1 4.178 (1.00$\times$) 6.042 (1.45$\times$) 10.446 (2.50$\times$) 4.237 (1.01$\times$) 6.094 (1.46$\times$) 10.415 (2.49$\times$)
SelfRegulationSCP2 3.295 (1.00$\times$) 4.67 (1.42$\times$) 9.391 (2.85$\times$) 3.269 (0.99$\times$) 4.629 (1.40$\times$) 9.376 (2.85$\times$)
SpokenArabicDigits 96.299 (1.00$\times$) 143.411 (1.49$\times$) 131.577 (1.37$\times$) 86.495 (0.90$\times$) 125.916 (1.31$\times$) 129.073 (1.34$\times$)
StandWalkJump 0.304 (1.00$\times$) 0.404 (1.33$\times$) 1.68 (5.53$\times$) 0.31 (1.02$\times$) 0.4 (1.32$\times$) 1.7 (5.59$\times$)
Avg. over non-spatial datasets 17.357 sec/epoch 1.46$\times$ 2.57$\times$ 0.99$\times$ 1.44$\times$ 2.57$\times$
In addition, to visually show the effect of differently regularised contrastive learning losses on representation, we apply t-SNE to compress the encoded representations into 3 dimensions, as plotted in Figure B.2 for the dataset Epilepsy, and Figure B.2 for RacketSports. The classes are indicated by colours. We use these two datasets because they are visualisation-friendly, with 4 classes and around 150 test samples.
Figure B.1: Encoded representations after training with different losses on the test set of Epilepsy.
Figure B.2: Encoded representations after training with different losses on the test set of RacketSports.
Appendix C Detailed results of macroscopic prediction with LSTM and GRU
This section provides additional tables presenting the evaluation of the final results using LSTM and GRU in macroscopic traffic prediction. Table C.1 shows the task-specific metrics and Table C.2 shows the metrics for global structure preservation.
Table C.1: Macroscopic traffic prediction evaluation with LSTM and GRU encoders.
@@ Appendix C Detailed results of macroscopic prediction with LSTM and GRU
EVar
No pretraining 3.244 6.402 6.400 82.161 3.552 7.228 7.228 77.248
TS2Vec 3.158 6.187 6.187 83.331 3.601 7.297 7.297 76.813
Topo-TS2Vec 3.139 6.155 6.154 83.505 3.491 7.006 7.006 78.626
GGeo-TS2Vec 3.101 5.974 5.973 84.461 3.466 6.948 6.947 78.984
SoftCLT 3.191 6.319 6.319 82.612 3.349 6.649 6.649 80.749
Topo-SoftCLT 3.192 6.271 6.270 82.877 3.367 6.661 6.660 80.682
GGeo-SoftCLT 3.128 6.121 6.121 83.683 3.302 6.495 6.495 81.630
Best improvement (%) 4.415 6.676 6.668 2.799 7.026 10.141 10.145 5.673
@@
Table C.2: Global structure preservation of LSTM and GRU encoders in macroscopic traffic prediction task.
@@
kNN Cont. Trust. MRRE dRMSE kNN Cont. Trust. MRRE dRMSE
No pretraining 0.161 0.858 0.899 0.097 0.427 0.151 0.870 0.902 0.091 0.424
TS2Vec 0.120 0.907 0.891 0.086 0.437 0.132 0.921 0.906 0.072 0.399
Topo-TS2Vec 0.122 0.899 0.888 0.089 0.452 0.134 0.914 0.904 0.076 0.429
GGeo-TS2Vec 0.134 0.912 0.905 0.078 0.427 0.138 0.916 0.913 0.071 0.426
SoftCLT 0.141 0.938 0.912 0.063 0.453 0.166 0.942 0.931 0.052 0.435
Topo-SoftCLT 0.145 0.932 0.917 0.061 0.444 0.175 0.951 0.934 0.046 0.471
GGeo-SoftCLT 0.147 0.940 0.918 0.059 0.477 0.177 0.948 0.935 0.046 0.456

### v4

@@ Acknowledgments
Appendix A Detailed results on UEA datasets
This section provides detailed comparisons of evaluation results for the used 28 datasets in the UEA archive. Tables A.1 and A.2 present the results of classification accuracy. Tables A.3 and A.4 present the training time for self-supervised representation learning.
Table A.1: Detailed evaluation of classification accuracy on spatial datasets in the UEA archive.
@@ Appendix A Detailed results on UEA datasets
Table A.2: Detailed evaluation of classification accuracy on non-spatial datasets in the UEA archive.
@@
Table A.3: Detailed representation training time per epoch (unit: s) on spatial datasets in the UEA archive.
@@
[…]
Table A.4: Detailed representation training time per epoch (unit: s) on non-spatial datasets in the UEA archive.
@@
[…]
In addition, to visually show the effect of differently regularised contrastive learning losses on representation, we apply t-SNE to compress the encoded representations into 3 dimensions, as plotted in Figure A.2 for the dataset Epilepsy, and Figure A.2 for RacketSports. The classes are indicated by colours. We use these two datasets because they are visualisation-friendly, with 4 classes and around 150 test samples.
Figure A.1: Encoded representations after training with different losses on the test set of Epilepsy.
Figure A.2: Encoded representations after training with different losses on the test set of RacketSports.
Appendix B Detailed results of macroscopic prediction with LSTM and GRU
This section provides additional tables and figures presenting the evaluation of the final results using LSTM and GRU in macroscopic traffic prediction. Table B.1 shows the task-specific metrics and Table B.2 shows the metrics for global structure preservation. Figures B.2 and B.2 show the latent representations encoded by LSTM and GRU models, respectively.
Table B.1: Macroscopic traffic prediction evaluation with LSTM and GRU encoders. Metrics are reported as mean $\pm$ standard deviation over 10 stratified folds of the test set.
@@ Appendix B Detailed results of macroscopic prediction with LSTM and GRU
EVar (↑)
No pretraining 3.244 $\pm$ 0.046 6.401 $\pm$ 0.102 6.399 $\pm$ 0.102 82.151 $\pm$ 0.512 3.552 $\pm$ 0.063 7.227 $\pm$ 0.119 7.227 $\pm$ 0.119 77.241 $\pm$ 0.457
TS2Vec 3.158 $\pm$ 0.045 6.187 $\pm$ 0.089 6.186 $\pm$ 0.089 83.322 $\pm$ 0.413 3.601 $\pm$ 0.062 7.296 $\pm$ 0.113 7.296 $\pm$ 0.113 76.805 $\pm$ 0.427
Topo-TS2Vec 3.139 $\pm$ 0.048 6.154 $\pm$ 0.107 6.153 $\pm$ 0.107 83.499 $\pm$ 0.423 3.491 $\pm$ 0.057 7.005 $\pm$ 0.108 7.005 $\pm$ 0.108 78.617 $\pm$ 0.442
GGeo-TS2Vec 3.101 $\pm$ 0.046 5.974 $\pm$ 0.089 5.973 $\pm$ 0.089 84.454 $\pm$ 0.354 3.466 $\pm$ 0.055 6.947 $\pm$ 0.105 6.946 $\pm$ 0.105 78.975 $\pm$ 0.417
SoftCLT 3.191 $\pm$ 0.049 6.319 $\pm$ 0.100 6.318 $\pm$ 0.100 82.604 $\pm$ 0.412 3.349 $\pm$ 0.051 6.649 $\pm$ 0.101 6.648 $\pm$ 0.101 80.741 $\pm$ 0.388
Topo-SoftCLT 3.192 $\pm$ 0.049 6.270 $\pm$ 0.101 6.270 $\pm$ 0.101 82.872 $\pm$ 0.317 3.367 $\pm$ 0.049 6.660 $\pm$ 0.098 6.659 $\pm$ 0.098 80.674 $\pm$ 0.398
GGeo-SoftCLT 3.128 $\pm$ 0.044 6.121 $\pm$ 0.094 6.120 $\pm$ 0.094 83.675 $\pm$ 0.400 3.302 $\pm$ 0.049 6.494 $\pm$ 0.097 6.494 $\pm$ 0.097 81.621 $\pm$ 0.430
Best improvement 4.415 6.675 6.667 2.803 7.026 10.139 10.143 5.671
@@
Table B.2: Global structure preservation of LSTM and GRU encoders in macroscopic traffic prediction task.
@@
kNN (↑) Trust. (↑) Cont. (↑) MRRE (↓) dRMSE (↓) kNN (↑) Trust. (↑) Cont. (↑) MRRE (↓) dRMSE (↓)
No pretraining 0.215 0.929 0.942 0.046 0.319 0.207 0.946 0.951 0.037 0.299
TS2Vec 0.173 0.959 0.955 0.035 0.318 0.197 0.963 0.957 0.031 0.269
Topo-TS2Vec 0.179 0.947 0.945 0.042 0.332 0.193 0.963 0.962 0.030 0.304
GGeo-TS2Vec 0.198 0.968 0.961 0.029 0.299 0.203 0.962 0.962 0.029 0.306
SoftCLT 0.186 0.954 0.948 0.038 0.334 0.214 0.958 0.957 0.031 0.319
Topo-SoftCLT 0.193 0.947 0.950 0.039 0.326 0.239 0.966 0.964 0.026 0.354
GGeo-SoftCLT 0.176 0.945 0.942 0.042 0.365 0.228 0.963 0.957 0.029 0.339
@@
Figure B.1: LSTM encoded representations after training with different losses on the test set of the Macroscopic traffic prediction task.
Figure B.2: GRU encoded representations after training with different losses on the test set of the Macroscopic traffic prediction task.